# Bayesian Prompt Flow Learning for Zero-Shot Anomaly Detection

### v1

@@ 3.5 Anomaly Map and Anomaly Score
Pixel-level anomaly map. We extract $L$-layer patch-level features and align them with the text embeddings. The anomaly map from the $i\text{-}th$ layer corresponding to text embeddings $\mathbf{F}_{b,r}^{t}$ is computed as:
$\mathbf{M}_{i,b,r}=softmax(Up(\widetilde{\mathbf{F}}^{I}_{i}\widetilde{\mathbf{F}}_{b,r}^{t\text{T}}))$ (11)
where $\widetilde{(\cdot)}$ denotes the $L_{2}$-normalization operation along the embedding dimension and $Up(\cdot)$ is the unsampling operation. The final result $\mathbf{M}\in\mathbb{R}^{H\times W}$ is obtained by averaging the anomaly maps derived from aligning the $L$-layer patch features with $B\times R$ text embeddings in the prompt banks:
$\mathbf{M}=\frac{1}{LBR}\sum_{i=1}^{L}\sum_{b=1}^{B}\sum_{r=1}^{R}\mathbf{M}_{i,b,r}$ (12)
Image-level anomaly score. For anomaly classification, the anomaly values come from both the text and image branches. The anomaly score in the text branch is computed as: $s_{text}=(\frac{1}{BR})\sum_{b}^{B}\sum_{r}^{R}softmax(\widetilde{\mathbf{x}}\widetilde{\mathbf{F}}_{b,r}^{t\text{T}})$, where $\mathbf{x}=\mathbf{x}_{{cls}}+\mathbf{x}_{{patch}}$ is the global image embedding. Note that $\mathbf{x}_{{cls}}$ refers to the global image features obtained from the class token of the vanilla image encoder, while $\mathbf{x}_{{patch}}$ comes from the fusion of fine-grained patch features. Specifically, the patch features $\mathbf{F}^{I}_{i}\in\mathbb{R}^{H\times W\times C}$ from different layers are first concatenated along the channel dimension. Then, global average pooling is applied along the spatial dimensions, followed by a linear layer that maps the results to $\mathbf{x}_{patch}$. The anomaly score from the image branch is derived from the maximum value of the anomaly map: $s_{img}=\max(\mathbf{M})$. The final image-level anomaly score is then expressed as: $s=s_{text}+s_{img}$.
@@ 3.6 Loss Function
During the training stage, a single Monte Carlo sampling is performed to enhance efficiency. Let the text prompts from $B$ samples be denoted as $[g_{b,1}^{n},g_{b,1}^{a}],b\in\{1,2,\cdots B\}$ with the corresponding text embeddings being $[\mathbf{t}_{b,1}^{n},\mathbf{t}_{b,1}^{a}],b\in\{1,2,\cdots B\}$. To enhance the diversity of learnable prompts in prompt banks, an orthogonal loss is designed on the text embeddings as follows:
@@
where the prompt flow loss $\mathcal{L}_{p}$ is computed using Equation (7). The first and second terms of $\mathcal{L}_{p}$ are distribution regularization terms, refining the initial distribution $q_{0}(\Phi_{0})$ into a simple prior $p(\Phi_{K})$ to effectively control the distribution of latent variables. This regularization ensures that the prompt flow module can learn an effective prompt distribution, mitigating overfitting while enhancing the generalization of the prompts. The third term, $E_{q_{0}(\Phi_{0})}[\log p(D|\Phi_{K})]$, maximizes the log-likelihood of the data, which is approximated by the sum of classification loss (text-image alignment) and segmentation loss (text-patch alignment). Specifically, our Bayes-PFL employs a cross-entropy loss for classification, along with the sum of Focal loss [35] and Dice loss [29] for segmentation. More details about the loss function can be found in Appendix A.1.
@@ 4.1 Experimental Setup
Datasets. To evaluate the ZSAD performance of the model, experiments on 15 real-world datasets from the industrial and medical domains are conducted. Specifically, we use seven commonly used industrial anomaly detection datasets, including MVTec-AD [4], VisA [47], BTAD [30], KSDD2 [6], RSDD [43], DAGM [12] and DTD-Synthetic [3]. Eight different datasets in the medical domain are used, including HeadCT [37], BrainMRI [37], and Br35H [16] for brain tumor classification, CVC-ColonDB [38], CVC-ClinicDB [5], Endo [17], and Kvasir [19] for colon polyp segmentation and ISIC [11] for skin cancer detection. Following [10, 46], we perform auxiliary training on one industrial dataset and directly infer on other industrial and medical datasets. Since the categories in VisA do not overlap with those in the other datasets, we use VisA as the auxiliary training set. To assess VisA itself, we fine-tune our model on the MVTec-AD dataset. More details about the datasets can be found in Appendix C.
@@
Implementation Details. The publicly available CLIP (ViT-L-14-336) pretrained by OpenAI [32] is adopted in this study. Following previous works [9, 46, 8], we standardize the resolution of input images to $518\times 518$ and select the 6th, 12th, 18th, 24th layers from the 24-layer image encoder to extract patch embeddings. The number of prompts $B$ in the prompt banks is set to 3 and flow length $K$ is set to 10 by default. The length of learnable context vectors $P$ and learnable state vectors $Q$ are both set to 5. The number of sampling iterations $R$ is set to 10 by default during the inference stage. We use the Adam [23] optimizer to train our Bayes-PFL for a total of 20 epochs, with an initial learning rate of 0.0001 and a batch size of 32. All experiments are conducted on a single NVIDIA 3090 GPU, with 5 random seeds selected. The experimental results are the averages of different object categories from the target dataset. More implementation details can be found in Appendix A.
@@ 4.2 Comparison with State-of-the-art methods
In this study, five SOTA methods are employed for comparison with our Bayes-PFL, including WinCLIP [18], APRIL-GAN [9], CLIP-AD [10], AnomalyCLIP [46], and AdaCLIP [8]. To ensure a fair comparison, we use the same backbone, input image resolution, and experimental settings (training on VisA and testing on other datasets) for all methods. More details regarding the comparison of these methods are shown in Appendix A.
@@
Quantitative Comparison. Table 1 presents the quantitative results of ZSAD across different datasets in both industrial and medical domains. Our Bayes-PFL achieves state-of-the-art performance on image-level and pixel-level metrics across nearly all datasets in comparison with other methods. For example, on the medical dataset ISIC, our Bayes-PFL outperforms the second-best approach by 3.8%, 9.5%, and 10.2% in pixel-level AUROC, PRO, and AP, respectively. The poor PRO performance of AdaCLIP is attributed to the high variance in its predicted anomaly scores and its limited ability to effectively detect large anomalous regions. In contrast, Bayes-PFL explores a larger prompt space by sampling from the distribution, enhancing the detection of large anomalous regions. Note that the anomaly detection results on the medical dataset are directly derived from the fine-tuned weights of the industrial dataset VisA, indicating that the feature spaces for anomalies in these two domains overlap to some extent.
@@
Qualitative Comparison. In Figure 5, we present visualizations of the anomaly maps generated for various datasets from both industrial and medical domains. It can be observed that our Bayes-PFL achieves more accuracy segmentation results and more complete anomaly region localization compared to other methods. Especially in categories from the medical domain (e.g., colon, skin), our method demonstrates a significant advantage. This indicates that the learned prompt distribution can cover a more comprehensive prompt space and capture the anomaly semantics in images from different novel categories.
@@
Performance comparison at different training epochs. In Figure 5, we compare APRIL-GAN [9] and AnomalyCLIP [46], which employ two different text prompt design strategies, with our Bayes-PFL across various training epochs. The pixel-level AP during testing is reported to evaluate ZSAD performance. For a fair comparison, only the input text prompt design strategies of these methods are adopted, while other components remain consistent with Bayes-PFL. These modified versions are referred to as APRIL-GAN+ and AnomalyCLIP+, respectively. Our Bayes-PFL demonstrates more stable and steadily improving AP throughout training, surpassing the other two methods and highlighting the advantages of our prompt flow learning approach. In contrast, APRIL-GAN+ and AnomalyCLIP+ exhibit performance declines after reaching their peak AP values, with the latter showing a more significant degradation. In Appendix B, we further explore this phenomenon and analyze the reasons behind it.
@@
Remark. Our Bayes-PFL outperforms other SOTA approaches for three main reasons: 1) The prompt flow module effectively regularizes the prompt space with ISD and IAD, thereby enhancing the model’s generalization performance on unseen categories; 2) The distribution sampling adaptively generates diverse text prompt to cover the prompt space. The ensemble of results from rich prompts enables better ZSAD performance in Bayes-PFL; 3) The RCA module facilitates cross-modal intersection and enhances the alignment capability of different modalities.
@@
In Bayes-PFL, ISD and IAD are modeled in the context and state word embedding spaces, respectively, referred to collectively as ISD-IAD. A natural question arises: why not use ISD-ISD or IAD-IAD instead? In Figure 5, we compare the absolute AP improvement achieved by using ISD-IAD over ISD-ISD and IAD-IAD across six industrial datasets. Higher values indicate a more pronounced advantage of our approach. It can be observed that all metrics are positive, indicating that ISD-IAD outperforms the other two distribution design strategies. This is because the text prompts need to learn unified normal and abnormal semantics across different categories, for which the static IAD is more suitable. The dynamic ISD incorporates visual features, aiding in the acquisition of rich contextual information (e.g., object background) in the context word embeddings.
@@ 4.4 Ablation
Influence of different components. Ablation studies on modules, loss functions, and classifications are conducted separately, as shown in Table 2. The module ablation study shows that ZSAD’s performance declines when the RCA module or any distribution type in the prompt flow module is removed. Notably, omitting ISD results in a greater performance drop than omitting IAD in all metrics, indicating that context distribution modeling conditioned on the image has a greater impact on generalization ability. Not using the orthogonal loss $\mathcal{L}_{ort}$ also leads to a decrease of approximately 1% in both image-level and pixel-level metrics. In the classification ablation study, the best performance is achieved when all components are included. This suggests that the results from the text and image branches complement each other, and integrating fine-grained patch features into the global image features $\mathbf{x}_{cls}$ from the vanilla CLIP model helps enhance classification performance.
Influence of the number of sampling iterations. The ZSAD performance on novel categories under different sampling iterations $R$ is presented in Table 3. Additionally, the standard deviation and average inference time per image (calculated from 50 images) are reported to assess the stability of the results and inference efficiency. As the number of sampling iterations increases, our method demonstrates improved zero-shot generalization performance and stability, but it also results in longer inference times. Consequently, we identified a trade-off at $R=10$, which is set as the default value in our experiments. More ablation studies on the hyperparameters can be found in Appendix B.
Influence of the ensemble mode. Given the text embeddings $\mathbf{t}{b,r}^{n}$ and $\mathbf{t}{b,r}^{a}$, we consider two ensemble modes: 1) Text ensemble, where text embeddings are averaged first and then aligned with image features; and 2) Image ensemble, where image features are aligned separately first, followed by averaging the anomaly scores or maps. Table 4 compares the ZSAD performance of these modes along with the average inference time per image. The results show that image ensemble outperforms text ensemble while maintaining comparable inference efficiency. As illustrated in Figure 6, image ensemble produces more complete detection results with higher confidence in anomalous regions. Therefore, we adopt image ensemble as the default approach for obtaining the final ZSAD results.

### v2

@@ 3.5 Anomaly Map and Anomaly Score
Pixel-level anomaly map. We extract $L$-layer patch-level features $\mathbf{F}^{I}_{l}$ and align them with the refined text embeddings $\mathbf{F}_{l,b,r}^{t},l=1,2,\cdots,L$. The anomaly map from the $l\text{-}th$ layer is computed as:
$\mathbf{M}_{l,b,r}=softmax(Up(\widetilde{\mathbf{F}}^{I}_{l}\widetilde{\mathbf{F}}_{l,b,r}^{t\text{T}}))$ (11)
where $\widetilde{(\cdot)}$ denotes the $L_{2}$-normalization operation along the embedding dimension and $Up(\cdot)$ is the unsampling operation. The final result $\mathbf{M}\in\mathbb{R}^{h\times w}$ is obtained by averaging the anomaly maps derived from aligning the $L$-layer patch features with $B\times R$ refined text embeddings:
$\mathbf{M}=\frac{1}{LBR}\sum_{l=1}^{L}\sum_{b=1}^{B}\sum_{r=1}^{R}\mathbf{M}_{l,b,r}$ (12)
Image-level anomaly score. For anomaly classification, the anomaly values come from both the text and image branches. The anomaly score in the text branch is computed as: $s_{text}=(\frac{1}{BR})\sum_{b}^{B}\sum_{r}^{R}softmax(\widetilde{\mathbf{x}}\widetilde{\mathbf{Z}}_{b,r}^{t\text{T}})$, where $\mathbf{x}=\mathbf{x}_{{cls}}+\mathbf{x}_{{patch}}$ is the global image embedding. Note that $\mathbf{x}_{{cls}}$ refers to the global image features obtained from the class token of the vanilla image encoder, while $\mathbf{x}_{{patch}}$ comes from the fusion of fine-grained patch features. Specifically, the patch features $\mathbf{F}^{I}_{l}\in\mathbb{R}^{H\times W\times C},l=1,2,\cdots,L$ from different layers are first concatenated along the channel dimension. Then, global average pooling is applied along the spatial dimensions, followed by a linear layer that maps the results to $\mathbf{x}_{patch}$. The anomaly score from the image branch is derived from the maximum value of the anomaly map: $s_{img}=\max(\mathbf{M})$. The final image-level anomaly score is then expressed as: $s=s_{text}+s_{img}$.
@@ 3.6 Loss Function
During the training stage, a single Monte Carlo sampling is performed to enhance efficiency. Let the text prompts from $B$ samples be denoted as $[g_{b,1}^{n},g_{b,1}^{a}],b\in\{1,2,\cdots B\}$, and their corresponding text embeddings as $[\mathbf{t}_{b,1}^{n},\mathbf{t}_{b,1}^{a}],b\in\{1,2,\cdots B\}$. To enhance the diversity of learnable prompts in prompt banks, an orthogonal loss is designed on the text embeddings as follows:
@@
where the prompt flow loss $\mathcal{L}_{p}$ is computed using Equation (7). The first two terms of $\mathcal{L}_{p}$ are distribution regularization terms that refine the initial distribution $q_{0}(\Phi_{0})$ into a simple prior $p(\Phi_{K})$, effectively controlling the distribution of latent variables. This regularization ensures that the prompt flow module can learn an effective prompt distribution, mitigating overfitting while enhancing the generalization of the prompts. The third term, $E_{q_{0}(\Phi_{0})}[\log p(D|\Phi_{K})]$, maximizes the log-likelihood of the data, which is approximated by the sum of classification loss (for text-image alignment) and segmentation loss (for text-patch alignment). Specifically, our Bayes-PFL employs a cross-entropy loss for classification, along with the sum of Focal loss [37] and Dice loss [30] for segmentation. More details about the loss function can be found in Appendix A.1.
@@ 4.1 Experimental Setup
Datasets. To evaluate the ZSAD performance of the model, experiments on 15 real-world datasets from the industrial and medical domains are conducted. Specifically, we utilize seven commonly used industrial anomaly detection datasets, including MVTec-AD [4], VisA [50], BTAD [31], KSDD2 [6], RSDD [46], DAGM [12] and DTD-Synthetic [3]. Eight different datasets in the medical domain are used, including HeadCT [39], BrainMRI [22], and Br35H [16] for brain tumor classification, CVC-ColonDB [40], CVC-ClinicDB [5], Endo [17], and Kvasir [19] for colon polyp segmentation and ISIC [11] for skin cancer detection. Following [10, 33], we perform auxiliary training on one industrial dataset and directly infer on other industrial and medical datasets. Since the categories in VisA do not overlap with those in the other datasets, we use VisA as the auxiliary training set. To assess VisA itself, we fine-tune our model on the MVTec-AD dataset. More details about the datasets can be found in Appendix C.
@@
Implementation Details. The publicly available CLIP (ViT-L-14-336) pretrained by OpenAI [34] is adopted in this study. Following previous works [9, 49, 8], we standardize the resolution of input images to $518\times 518$ and select the 6th, 12th, 18th, 24th layers from the 24-layer image encoder to extract patch embeddings. The number of prompts $B$ in the prompt banks is set to 3 and flow length $K$ is set to 10 by default. The length of learnable context vectors $P$ and learnable state vectors $Q$ are both set to 5. During the auxiliary training and inference stages, the number of Monte Carlo sampling interations $R$ is set to 1 and 10, respectively. We use the Adam [24] optimizer to train our Bayes-PFL for a total of 20 epochs, with an initial learning rate of 0.0001 and a batch size of 32. All experiments are conducted on a single NVIDIA 3090 GPU, with 5 random seeds selected. The experimental results are the averages of different object categories from the target dataset. More implementation details can be found in Appendix A.1.
@@ 4.2 Comparison with State-of-the-art methods
In this study, five SOTA methods are employed for comparison with our Bayes-PFL, including WinCLIP [18], APRIL-GAN [9], CLIP-AD [10], AnomalyCLIP [49], and AdaCLIP [8]. To ensure a fair comparison, we use the same backbone, input image resolution, and experimental settings (training on VisA and testing on other datasets) for all methods. More details regarding the comparison of these methods are shown in Appendix A.2.
@@
Quantitative Comparison. Table 1 presents the quantitative results of ZSAD across different datasets in both industrial and medical domains. Our Bayes-PFL achieves state-of-the-art performance on image-level and pixel-level metrics across nearly all datasets in comparison with other methods. For example, on the medical dataset ISIC, our Bayes-PFL outperforms the second-best approach by 3.8%, 9.5%, and 10.2% in pixel-level AUROC, PRO, and AP, respectively. The poor PRO performance of AdaCLIP is attributed to the high variance in its predicted anomaly scores and its limited ability to effectively detect large anomalous regions. In contrast, Bayes-PFL explores a broader prompt space by sampling from the distribution, enhancing the detection of large anomalous regions. Note that the anomaly detection results on the medical dataset are directly derived from the fine-tuned weights of the industrial dataset VisA, indicating that the feature spaces for anomalies in these two domains overlap to some extent.
@@
Qualitative Comparison. Figure 5 visualizes anomaly maps from both industrial and medical datasets. It can be observed that our Bayes-PFL achieves more accurate segmentation results and more complete anomaly region localization compared to other methods. Especially in categories from the medical domain (e.g., colon, skin), our method demonstrates a significant advantage. These results indicate that the learned prompt distribution can cover a more comprehensive prompt space and capture the anomaly semantics in images from different novel categories.
@@
Performance comparison at different training epochs. In Figure 5, we compare APRIL-GAN [9] and AnomalyCLIP [49], which employ two different text prompt design strategies, with our Bayes-PFL across various training epochs. The pixel-level AP during testing is reported to evaluate ZSAD performance. For a fair comparison, we modify these methods by replacing only their text prompt designs while keeping all other components identical to Bayes-PFL. These modified versions are referred to as APRIL-GAN+ and AnomalyCLIP+, respectively. Our Bayes-PFL achieves stable and consistently improving AP throughout training, surpassing the other two methods and highlighting the advantages of our prompt flow learning approach. In contrast, APRIL-GAN+ and AnomalyCLIP+ exhibit performance declines after reaching their peak AP values, with the latter showing a more significant degradation. In Appendix B.4, we further explore this phenomenon and analyze the reasons behind it.
@@
Remark. Our Bayes-PFL outperforms other SOTA approaches for three main reasons: 1) The prompt flow module effectively regularizes the prompt space with ISD and IAD, thereby enhancing the model’s generalization performance on unseen categories; 2) The distribution sampling adaptively generates diverse text prompts to cover the prompt space. The ensemble of results from rich prompts enables better ZSAD performance in Bayes-PFL; 3) The RCA module facilitates cross-modal intersection and enhances the alignment capability of different modalities.
@@
In Bayes-PFL, ISD and IAD are modeled in the context and state word embedding spaces, respectively, referred to collectively as ISD-IAD. A natural question arises: why not use ISD-ISD or IAD-IAD instead? In Figure 5, we compare the absolute AP improvement achieved by using ISD-IAD over ISD-ISD and IAD-IAD across six industrial datasets. Higher values indicate a more pronounced advantage of our approach. It can be observed that all metrics are positive, indicating that ISD-IAD outperforms the other two distribution design strategies. This is because the text prompts need to learn unified normal and abnormal semantics across different categories, making the static IAD more suitable for capturing category-invariant state representations. The dynamic ISD incorporates visual features, aiding in the acquisition of rich contextual information (e.g., object background) in the context word embeddings.
@@ 4.4 Ablation
Influence of different components. Ablation studies on modules, loss functions, and classifications are conducted separately, as summarized in Table 2. The module ablation study shows that ZSAD performance declines when the RCA module or any distribution type in the prompt flow module is removed. Notably, omitting ISD results in a greater performance drop than omitting IAD in all metrics, indicating that context distribution modeling conditioned on the image has a greater impact on generalization ability. Not using the orthogonal loss $\mathcal{L}_{ort}$ also leads to a decrease of approximately 1% in both image-level and pixel-level metrics. In the classification ablation study, the best performance is achieved when all components are included. This suggests that the results from the text and image branches complement each other, and integrating fine-grained patch features into the global image features $\mathbf{x}_{cls}$ from the vanilla CLIP model helps enhance classification performance.
Influence of the number of sampling iterations. The ZSAD performance on novel categories under different sampling iterations $R$ is presented in Table 3. Additionally, the standard deviation and average inference time per image (calculated from 50 images) are reported to assess the stability of the results and inference efficiency. As the number of sampling iterations increases, our method demonstrates improved zero-shot generalization performance and stability, but at the cost of longer inference times. Consequently, we identified a trade-off at $R=10$, which is set as the default value in our experiments. More ablation studies on the hyperparameters can be found in Appendix B.
Influence of the ensemble mode. Given the text embeddings $\mathbf{t}_{b,r}^{n}$ and $\mathbf{t}_{b,r}^{a}$, we consider two ensemble modes: 1) Text ensemble, where text embeddings are averaged first and then aligned with image features; and 2) Image ensemble, where image features are aligned separately first, followed by averaging the anomaly scores or maps. Table 4 compares the ZSAD performance of these modes along with the average inference time per image. The results show that image ensemble outperforms text ensemble while maintaining comparable inference efficiency. As illustrated in Figure 6, image ensemble produces more complete detection results with higher confidence in anomalous regions. Therefore, we adopt image ensemble as the default approach for generating the final ZSAD results.
@@ 5 Conclusion
Acknowledgement
This work is supported in part by the National Science and Technology Major Project of China under Grant 2022ZD0119402; in part by the National Natural Science Foundation of China under Grant No 62373350 and 62371179; in part by the Youth Innovation Promotion Association CAS (2023145); in part by the Beijing Nova Program 20240484687.